A compressed air vessel has an inside diameter of 2 m and a plate thickness of 12,5 mm. If the tensile stress induced in the plating is to be limited to 100 MPa, calculate for both the longitudinal and circumferential seams the pressure necessary to induce this stress. Take the longitudinal and circumferential joint efficiencies as 80 % and 50 % respectively. What would the actual allowable pressure be?
1
Expert's answer
2020-12-07T03:48:49-0500

Diameter of vessel is 2m, and thickness of plate is 12.5 mm= 12.5 "\times 10 ^{-3} m"

Tensile stress= 100 MPa, Longitudinal efficiency= 80 % and

circumferential efficiency=50 %

Longitudinal stress= "\sigma _l= \frac{pd}{2t \eta_c}=\frac{100\times 2}{2\times 12.5\times 10^{-3}\times 0.5}= 1.6" Mpa

and circumferential stress

"\sigma _c= \frac{pd}{2t \eta_l}=\frac{100\times 2}{2\times 12.5\times 10^{-3}\times 0.8}= 1.0" Mpa

So allowable stress is less in both which is

1Mpa